Find the equivelent expression without perenthisis
-(a-68)=

HELP!!

Answer by nycsharkman(136)

Find the equivalent expression without perentheses.
-(a-68)=
Just apply the distributive rule.
-1 times a = -a
-1 times -68 = 68
Answer: -a + 68
There is always an invisible 1 in front of the parentheses in math. This time it happens to be -1.
So, we mutliply everything inside the parentheses by -1 and
the answer is -a + 68.